1) She wants the patio to have the most area possible using one of the given shapes. Which option gives her the most space?
2) The stone used for the edging of each shape is $12 per linear foot. Calculate the cost of the edging of each shape and explain which option is least expensive.
a) Square with a length of 9 3/4 feet.
b) Circle with a diameter of 10.95 feet.
Need the answer, ASAP.
Thank you!

Answer by Alan3354(69443)

1) She wants the patio to have the most area possible using one of the given shapes. Which option gives her the most space?
-----
a) Square with a length of 9 3/4 feet
If you mean side length, area = 9.75^2 = 95.0625 sq ft
----
b) Circle with a diameter of 10.95 feet.
Area = pi*r^2 = pi*5.475^2
=~ 94.171 sq ft
===========================
2) The stone used for the edging of each shape is $12 per linear foot. Calculate the cost of the edging of each shape and explain which option is least expensive.
---
Square:
Perimeter = 4*9.75 = 39 ft
----
For the circle:
C = pi*d =~ 34.4 ft